Grade upto college level IIT JEE Entrance Exam

E, m, J and G denote energy, mass, angular momentum and gravitational constant respectively, then the dimensions of (EJ^2)/(m^5 G^2 ) are that of
(a) Angle
(b) Length
(c) Mass
(d) Time

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Simran Bhatia
12 Years ago

(a)

    (EJ^2)/(M^5 G^2 )=[ML^2 T^(-2) ][ML^2 T^(-1) ]/([M^5 ] [M^(-1) L^3 T^(-2) ]^2 )=(M^3 L^6 T^(-4))/(M^3 L^6 T^(-4) )